When 15g of Hydrogen is combined with 12g of Oxygen, 20g of water formwith extra Oxygen left over. What is the mass of the excess Oxygen?
1
Expert's answer
2020-11-17T10:34:55-0500

2H2 + O2= 2H2O

The limiting reagent here is O2, therefore H2 and not O2 will be in excess

No of moles of H2 = 15/4 = 3.75 moles

No of moles of O2 =12/32= 0.375

Only 2×0.375 moles of H2 will react with 0.375mole of O2

No of moles of H2 that will react is 0.75

No of excess mol= 3.75-0.75 = 3moles of H2

1 mole of H2 is 2g

3 moles of H2 = 3×2= 6g of H2 in excess